Huilan is 6 years older that Thomas. The sum of their ages is 72. What is Thomas's age?
Answer by Earlsdon(6294)

First, assign the variables.
Let H = Huilan's age
T = Thomas's age.
H = T + 6
H + T = 72 Subsitute H from the first equation into the second equation.
(T + 6) + T = 72 Solve for T.
2T + 6 = 72 Subtract 6 from both sides.
2T = 66 Divide both sides by 2.
T = 33 Thomas's age is 33 years.
H = T + 6 = 33 + 6 = 39 Huilan's age is 39 years.
Check:
T+6 = 33+6 = 39
T+H = 33+39 = 72